a particle is projected vertically upward from a pont A on the ground . It takes t1time to reavh a point B but it still continues to mov up . If it takes further t2 time to reach the ground from point B then the height of point B from the Ground is
 
  1. 1/2g (t1+t2)2
  2. gt1t2
  3. 1/8g(t1+t2)2
  4. 1/2gt1+t2

Grade:11

2 Answers

Arun
4 years ago
Time taken for the particle to reach highest point is t1+t22t1+t22
(v=u−gt)(v=u−gt)
Therefore initial velocity=u=g(t1+t2/2)=u=g(t1+t2/2)
Therefore heigh of B from ground is
h=ut+1/2h=ut+1/2at2at2
h=ut1−1/2h=ut1−1/2gt12gt12
=g(t1+t2/2)=g(t1+t2/2)t1−12t1−1/2gt21gt12
h=1/2h=1/2gt1t2
Vikas TU
4 years ago
Dear student 
Kindly provide the image of question , 
Beciuse solution is not matching with the ans. 
We will help you 
Good Luck